# Enhancement of superconductivity outside Abrikosov vortex core in a tightly bound Cooper pair superconductor

### v2

@@ II Feynman’s phenomenological theory
[…]
@@
×$\nabla\times\mathbf{v}=-\frac{e}{mc}\nabla\times\mathbf{A}=-\frac{e}{mc}\mathbf{H}$ (14)
@@
$-\hbar\left(\frac{\partial\theta}{\partial t}+\frac{e}{\hbar}\phi\right)=\frac{m\mathbf{v}^{2}}{2}+V^{\prime}(n)-\frac{\hbar^{2}}{2m}\frac{\nabla^{2}\sqrt{n}}{\sqrt{n}}.$ (15)
@@
×$\nabla\times\textbf{H}=\frac{1}{c}\frac{\partial\textbf{E}}{\partial t}+\frac{4\pi}{c}en\mathbf{v}.$ (17)
@@ III Stationary field equations
×$\nabla\times\textbf{H}=\frac{4\pi}{c}en\mathbf{v}.$ (22)
@@ IV Abrikosov flux line
[…]
@@ IV.1 Field configuration
[…]
@@ IV.1.2 Large distance behavior: weak-screening limit
which has a form of the unscreened potential of a charged line: electrostatic screening is ineffective at small distances. This result is consistent with the expectation that (in the original physical units) the region of the vortex core $\rho\lesssim\xi$ is positively charged with linear charge density of the order $-en_{0}\xi^{2}$.
@@ IV.1.3 Large distance behavior: strong-screening limit
Comparing the first and second terms and ignoring the integer $l$, we deduce that the strong screening approximation and thus the limiting solution (45) and (48) apply up to distances of the order $\sqrt{\gamma}$. Consequently, the net (dimensionless) charge (per unit length) accumulated at distances $\rho\gtrsim\sqrt{\gamma}$ can be estimated as $-\gamma$. It is compensated by the net positive charge of the same order $\gamma$ for $\rho\lesssim\sqrt{\gamma}$ to ensure global neutrality. Therefore, the length scale $\sqrt{\gamma}$ has a meaning of the size of the vortex core. Since $\gamma\ll 1$, the size of the vortex core is substantially smaller than that in the weak screening regime. In physical units, the size of the vortex core in the strong screening regime is $\xi\sqrt{\gamma}=\xi\sqrt{\delta/\xi}=\sqrt{\delta\xi}$, whereas it is simply $\xi$ when electrostatic screening is weak (or electric effects are disregarded).
@@
[…]
@@ IV.2 Numerical results
The problem of field configurations due the Abrikosov flux line accounting for electrostatic effects was also studied numerically. Since $\kappa\gg 1$ and most significant deviations from Abrikosov’s original predictions Abrikosov take place at small distances from the vortex axis, $\rho\lesssim\kappa$, we carried out a detailed study of the purely electric version of the problem, $\kappa=\infty$. Then, according to Ampère’s law (33), we only need to solve the $F=1$ case of Eq.(IV.1) together with the Poisson equation (32).

### v3

@@ II Feynman’s phenomenological theory
$S=\int\mathcal{L}\left(q,\frac{\partial q}{\partial t},\nabla q,\mathbf{r},t\right)dtd^{3}x$ (1)
@@
$\nabla\times\mathbf{v}=-\frac{e}{mc}\nabla\times\mathbf{A}=-\frac{e}{mc}\mathbf{H}$ (14)
@@
$\hbar\left(\frac{\partial\theta}{\partial t}+\frac{e}{\hbar}\phi\right)=\frac{m\mathbf{v}^{2}}{2}+V^{\prime}(n)-\frac{\hbar^{2}}{2m}\frac{\nabla^{2}\sqrt{n}}{\sqrt{n}}.$ (15)
@@
$\nabla\times\textbf{H}=\frac{1}{c}\frac{\partial\textbf{E}}{\partial t}+\frac{4\pi}{c}en\mathbf{v}.$ (17)
@@ III Stationary field equations
$\nabla\times\textbf{H}=\frac{4\pi}{c}en\mathbf{v}.$ (22)
@@ IV Abrikosov flux line
$\int\mathbf{H}\cdot d\mathbf{s}=\oint\mathbf{A}\cdot d\mathbf{l}=\frac{\hbar c}{e}\oint\nabla\theta\cdot d\mathbf{l}=\frac{2\pi\hbar c}{e}l$ (26)
@@ IV.1 Field configuration
$\oint\mathbf{v}\cdot d\mathbf{l}=\frac{\hbar}{m}\oint\nabla\theta\cdot d\mathbf{l}=\frac{2\pi\hbar}{m}l.$ (37)
@@ IV.1.2 Large distance behavior: weak-screening limit
which has a form of the unscreened potential of a charged line: electrostatic screening is ineffective at small distances. This result is consistent with the expectation that (in the original physical units) the region of the vortex core $\rho\lesssim\xi$ is positively charged with linear charge density of the order $en_{0}\xi^{2}$.
@@ IV.1.3 Large distance behavior: strong-screening limit
Comparing the first and second terms and ignoring the integer $l$, we deduce that the strong screening approximation and thus the limiting solution (45) and (48) apply up to distances of the order $\sqrt{\gamma}$. Consequently, the net (dimensionless) charge (per unit length) accumulated at distances $\rho\gtrsim\sqrt{\gamma}$ can be estimated as $\gamma$. It is compensated by the net positive charge of the same order $\gamma$ for $\rho\lesssim\sqrt{\gamma}$ to ensure global neutrality. Therefore, the length scale $\sqrt{\gamma}$ has a meaning of the size of the vortex core. Since $\gamma\ll 1$, the size of the vortex core is substantially smaller than that in the weak screening regime. In physical units, the size of the vortex core in the strong screening regime is $\xi\sqrt{\gamma}=\xi\sqrt{\delta/\xi}=\sqrt{\delta\xi}$, whereas it is simply $\xi$ when electrostatic screening is weak (or electric effects are disregarded).
@@
$\varepsilon_{l}$ $=$ $\int\frac{mn\mathbf{v}^{2}}{2}dxdy=\frac{\pi\hbar^{2}}{m}l^{2}\int_{0}^{\infty}\frac{n(\rho)F^{2}(\rho)}{\rho}d\rho$ (51)
@@ IV.2 Numerical results
The problem of field configurations due the Abrikosov flux line accounting for electrostatic effects was also studied numerically. Since $\kappa\gg 1$ and most significant deviations from Abrikosov’s original predictions Abrikosov take place at small distances from the vortex axis, $\rho\lesssim\kappa$, we carried out a detailed study of the purely electric version of the problem, $\kappa=\infty$. Then, according to Ampère’s law (33), we only need to solve the $F=1$ case of Eq.(IV.1) together with the Poisson equation (32); a code is available online Mia .